Grade 12th passElectromagnetic Induction

how does current changes with time in LR circuit across the inductor.

Vaibhav Tamta
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
In a simple L-R circuit,
 
If apply KVL we get,
 
L*di/dt + iR = 0
integeratig both sides via variable separable method,
we get,
di/i = -R/Ldt
integerate both sides now,
 
log(i) = -(R/L)t
i = e^(-R/L)*t
 
is the required answer.